OpenStudy (asylum15):

Q. Find the area under the curve of:

OpenStudy (asylum15):

y (x) = x(1-x) between x = 0.5 and x = 1

OpenStudy (anonymous):

\[\int\limits_{0.5}^{1}x(1-x)\] It will be an integral to solve ...

OpenStudy (asylum15):

Taking (1-x) as U?

OpenStudy (amistre64):

...just multiply x by 1-x and work the poly

OpenStudy (asylum15):

Could you elaborate mate? :)

OpenStudy (amistre64):

what is x(1-x) ? this is just simple algebra

OpenStudy (asylum15):

x - x^2

OpenStudy (amistre64):

good, then its just some basic power integrations

OpenStudy (asylum15):

This Q is on my engineering technological maths exam, so I expected it to be harder than it looks :D

OpenStudy (amistre64):

but why? its just an engineering book lol

OpenStudy (asylum15):

Do you mean integrate the x-x^2 yeah?

OpenStudy (asylum15):

x = x^2 / 2 eg

OpenStudy (amistre64):

yep, add1 and divide is the basic power rule

OpenStudy (asylum15):

Then fill it the bounds?

OpenStudy (amistre64):

and recall:\[\int_{a}^{b}f(x)~dx=F(b)-F(a)\]

OpenStudy (asylum15):

Yes

OpenStudy (amistre64):

so, what do you get, so we can chk if its right :)

OpenStudy (asylum15):

\[[\frac{ x^2 }{ x } - \frac{ x^3 }{ x } ]\] for (b) minus the same equation for (a)

OpenStudy (amistre64):

\[f=x-x^2\] \[F=\frac12x^2-\frac13x^3\] F(b): F(a): \[(\frac121^2-\frac131^3)-(\frac12.5^2-\frac13.5^3)\] \[(\frac16)-(\frac18-\frac1{24})\] etc

OpenStudy (asylum15):

Could I leave the equation as x^2 /x - x^3/3 and fill in for the bounds

OpenStudy (amistre64):

x^2 / 2 not x^2/x and what do you mean by "fill in for the bounds"? determine F(b) , then determine F(a) using the function F(x) that was integrated then subtract F(a) from F(b)

OpenStudy (asylum15):

I'm on the same page. Thank you Amistre! May I ask you one other thing quickly? Will help me hugely for my exam.

OpenStudy (amistre64):

sure, as long as i got time to study for my sociology final ...

OpenStudy (asylum15):

LOL! Really quick

OpenStudy (asylum15):

Q. Find the mean value of the function f (t) = \[10e^{-2t}\] between t = 0 and t = 2.5

OpenStudy (amistre64):

 if i recall correctly, it is the value of f(t) that would create the same area but with a level line